# Improving Low-Resource Knowledge Tracing Tasks by Supervised Pre-training and Importance Mechanism Fine-tuning

### v2

@@ 5.3 Baselines
DKT (Piech et al., 2015) Sequential 0.7228 0.7031 0.8105 0.8554 0.8173 0.6665
DKVMN (Zhang et al., 2017) Memory 0.7196 0.7022 0.8026 0.8547 0.8155 0.6392
DKT+ (Yeung and Yeung, 2018) Sequential 0.7243 0.7046 0.8085 0.8554 0.8179 0.6668
DKT-F (Nagatani et al., 2019) Sequential - 0.7076 0.8092 0.8540 0.8209 0.6666
KQN (Lee and Yeung, 2019) Sequential 0.7214 0.7028 0.8022 0.8540 0.8154 0.6665
SKVMN (Abdelrahman and Wang, 2019) Memory 0.7156 0.6885 0.7837 0.8406 0.8071 0.6572
DeepIRT (Yeung, 2019) Memory 0.7196 0.7020 0.8029 0.8542 0.8152 0.6559
GKT (Nakagawa et al., 2019) Others 0.7179 0.7053 0.8078 0.8555 0.8139 0.6672
SAKT (Pandey and Karypis, 2019) Attention 0.7031 0.6873 0.7948 0.8460 0.8121 0.6519
SAINT (Choi et al., 2020) Attention 0.6977 0.7187 0.7770 0.8455 0.8177 0.6624
AKT (Ghosh et al., 2020) Attention 0.7385 0.7320 0.8138 0.8594 0.8275 0.6888
ATKT (Guo et al., 2021) Others 0.7206 0.7012 0.7989 0.8510 0.8143 0.6642
HawkesKT (Wang et al., 2021) Others 0.7046 0.7110 0.8108 0.8563 0.8191 0.7076
LPKT (Shen et al., 2021) Sequential 0.7355 0.7309 0.8154 0.8547 0.8264 0.7243
AT-DKT (Liu et al., 2023) Sequential 0.7250 0.7146 0.8144 0.8560 0.8195 0.6684
simpleKT (Liu et al., 2023) Attention 0.7320 0.7328 0.8083 0.8579 0.8240 0.6624
sparseKT (Huang et al., 2023) Attention 0.7282 0.7322 0.8017 0.8569 0.8234 0.6643

### v3

@@ 5.3 Baselines
DKT [16] Sequential 0.7228 0.7031 0.8105 0.8554 0.8173 0.6665
DKVMN [7] Memory 0.7196 0.7022 0.8026 0.8547 0.8155 0.6392
DKT+ [17] Sequential 0.7243 0.7046 0.8085 0.8554 0.8179 0.6668
DKT-F [8] Sequential - 0.7076 0.8092 0.8540 0.8209 0.6666
KQN [18] Sequential 0.7214 0.7028 0.8022 0.8540 0.8154 0.6665
SKVMN [55] Memory 0.7156 0.6885 0.7837 0.8406 0.8071 0.6572
DeepIRT [56] Memory 0.7196 0.7020 0.8029 0.8542 0.8152 0.6559
GKT [26] Others 0.7179 0.7053 0.8078 0.8555 0.8139 0.6672
SAKT [19] Attention 0.7031 0.6873 0.7948 0.8460 0.8121 0.6519
SAINT [36] Attention 0.6977 0.7187 0.7770 0.8455 0.8177 0.6624
AKT [20] Attention 0.7385 0.7320 0.8138 0.8594 0.8275 0.6888
ATKT [10] Others 0.7206 0.7012 0.7989 0.8510 0.8143 0.6642
HawkesKT [58] Others 0.7046 0.7110 0.8108 0.8563 0.8191 0.7076
LPKT [53] Sequential 0.7355 0.7309 0.8154 0.8547 0.8264 0.7243
AT-DKT [54] Sequential 0.7250 0.7146 0.8144 0.8560 0.8195 0.6684
simpleKT [24] Attention 0.7320 0.7328 0.8083 0.8579 0.8240 0.6624
sparseKT [57] Attention 0.7282 0.7322 0.8017 0.8569 0.8234 0.6643